OpenStudy (anonymous):

A bag contains 5 blue marbles, 2 black marbles, and 3 red marbles. A marble is randomly drawn from the bag. The probability of not drawing a black marble is? The probability of drawing a red marble is?

OpenStudy (anonymous):

@Kitten_is_back

OpenStudy (anonymous):

@Error1603

OpenStudy (anonymous):

I need an answer quick

OpenStudy (error1603):

10 marbles total, 2 red 2/10 x 2/10 = 4/100= 1/25 2/10 x 1/9 = 2/90 = 1/45 ( marble not replaced so you have one less the second time)

OpenStudy (anonymous):

so what is the probability to not draw a black marble

OpenStudy (error1603):

Yes

OpenStudy (anonymous):

what is the answer